# Enhancing Variable Selection in Large-scale Logistic Regression: Leveraging Manual Labeling with Beneficial Noise

### v1

@@ V Numerical Experiments
In this section, we validate that manual labeling can enhance the variable selection ability of PLR on datasets in Table I. These datasets is available at https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/. Due to space limitations, we only present the numerical results of the w8a dataset with the largest data size in the table. The numerical results of other datasets, artificially synthesized data, comparison of algorithm efficiency can be found in the online appendix.
TABLE I: Basic Information of Datasets
Dataset Features (d) Samples (n)
ijcmn1 22 49,990
phishing 68 11,055
a9a 123 32,561
w8a 300 49,749
@@
We randomly partitioned the original dataset comprising 49,749 observations into a training set of 40,000 observations and a test set of 9,749 observations. Table II presents the simulated relative efficiency of PLR. Here, $\bm{\beta}^{*}$ is an estimation using the entire sample data. The simulated ARE values tend to increase as $\alpha_{0}$ and $m$ increase, which is consistent with the conclusion of Theorem 2. It is observed that the simulated relative efficiencies are lower than their corresponding theoretical values. This discrepancy might stem from two factors. Firstly, there could be an underestimation of the conditional error rate $\text{err}(\hat{\bm{\beta}};\bm{\beta}^{*})$ in (18). Secondly, the true conditional error rate $\text{err}({\bm{\beta}}^{*};\bm{\beta}^{*})$ may not be precisely estimated using the limited test set. In the simulations presented in online appendix, we were able to compute the exact $\text{err}(\hat{\bm{\beta}};\bm{\beta}^{*})$ and $\text{err}({\bm{\beta}}^{*};\bm{\beta}^{*})$. This ability can account for the better agreement between the simulated relative efficiency in appendix and the asymptotic relative efficiency observed in this section.
@@
TABLE II: The Asymptotic Relative Efficiency (ARE) On the w8a Dataset. The Theoretical ARE of PLR, as stated in Theorem 2, is provided in parentheses.
@@
VI Conclusion and Future Work
This paper theoretically proves that when label noise is only related to classification difficulty, it is beneficial for variable selection in binary classification PLR. This benefit will be reflected in more accurate estimates of important coefficients compared to using truth labels. In addition, we also propose a new parallel algorithm for solving the large-scale PLR with manually generated labels. This parallel algorithm is insensitive to data partitioning, meaning that the solution of the algorithm remains unchanged regardless of different data partitions.
Future research can be extended in several directions: expanding to multi-class classification to see if the beneficial effect of label noise persists; optimizing the parallel algorithm for better computational efficiency and scalability using techniques like GPU acceleration; incorporating temporal and spatial information into the PLR model and algorithm to handle data with such characteristics.
Acknowledgements
We express our sincere gratitude to Professor Bingsheng He for engaging in invaluable discussions with us. The research of Wu was supported by the Scientific and Technological Research Program of Chongqing Municipal Education Commission [Grant Numbers KJQN202302003].
Appendix A Proof
A-A Proof of Throrem 1
@@
$\Gamma_{n}(\mathbf{u})$ $=\sum_{i=1}^{n}\left(-S_{i}\left(\bm{x}_{i}^{\top}\left(\bm{\beta}^{*}+\frac{\bm{u}}{\sqrt{n}}\right)\right)\right.$
$+\phi\left(\bm{x}_{i}^{\top}\left(\bm{\beta}^{*}+\frac{\bm{u}}{\sqrt{n}}\right)\right)+\lambda\sum_{j=1}^{d}w_{j}\left|\beta_{j}^{*}+\frac{u_{j}}{\sqrt{n}}\right|.$
Let $\tilde{\bm{u}}^{(n)}=\arg\min\limits_{\bm{u}}\Gamma_{n}(\bm{u})$; then $\tilde{\bm{u}}^{(n)}=\sqrt{n}(\tilde{\bm{\beta}}-\bm{\beta}^{*})$, where $\tilde{\bm{\beta}}$ is the PLR estimator with manual labels in (17).
@@
$\mathbb{E}_{S_{i},\bm{x}_{i}}\left(\left[S_{i}-\phi^{\prime}\left(\bm{x}_{i}^{\top}\bm{\beta}^{*}\right)\right]\bm{x}_{i}^{\top}\right)=\bm{0}$ (40)
@@
$\text{Var}_{S_{i},\bm{x}_{i}}\left(\left[S_{i}-\phi^{\prime}\left(\bm{x}_{i}^{\top}\bm{\beta}^{*}\right)\right]\bm{x}_{i}^{\top}\right)=\frac{m+\alpha_{0}}{1+\alpha_{0}}mI(\bm{\beta}^{*}).$ (41)
@@
$\text{Var}_{S_{i},\bm{x}_{i}}\left(\left[S_{i}-\phi^{\prime}\left(\bm{x}_{i}^{\top}\bm{\beta}^{*}\right)\right]\left(\bm{x}_{i}^{\top}\bm{u}\right)\right)$
$=$ $\bm{u}^{\top}\frac{m+\alpha_{0}}{1+\alpha_{0}}mI(\bm{\beta}^{*})\bm{u}.$
@@
[…]
For the second term $A_{2}^{(n)}$, from (4), we observe that
[…]
@@
$\frac{\lambda}{\sqrt{n}}{w}_{j}\sqrt{n}\left(\left|\beta_{j}^{*}+\frac{u_{j}}{\sqrt{n}}\right|-\left|\beta_{j}^{*}\right|\right)$
[…]
@@
$6\sqrt{n}A_{4}^{(n)}\leq$ $\sum_{i=1}^{n}\frac{1}{n}M(\bm{x}_{i}^{\top})|\bm{x}_{i}^{\top}\bm{u}|^{3}$
[…]
@@
[…]
@@
$\sum_{i=1}^{n}x_{ij^{\prime}}\left(S_{i}-\phi^{\prime}\left(\bm{x}_{i}^{\top}\tilde{\bm{\beta}}\right)\right)=\lambda w_{j};$ (48)
@@
where $\tilde{\bm{\beta}}^{*}$ is between $\tilde{\bm{\beta}}$ and ${\bm{\beta}}^{*}$. From (40) and (41), we know that
[…]
It follows from (A-A) that
[…]
By considering the left and right sides of the (48), it can be concluded that $\Pr(j^{\prime}\in\tilde{\mathcal{A}})\to 0$. This completes the proof of Theorem 1.
A-B Proof of Throrem 2
From the conclusions of Section II-A2 and Theorem 1, it can be known that when $n\to\infty$, $\bm{\beta}^{*}=({\bm{\beta}^{*}_{\mathcal{A}}}^{\top},\bm{0}^{\top})^{\top}$, $\hat{\bm{\beta}}=(\hat{\bm{\beta}}_{\mathcal{A}}^{\top},\bm{0}^{\top})^{\top}$ and $\tilde{\bm{\beta}}=(\tilde{\bm{\beta}}_{\mathcal{A}}^{\top},\bm{0}^{\top})^{\top}$. This also means that $\text{err}(\hat{\bm{\beta}};\bm{\beta}^{*})=\sum\limits_{i=1}^{n}\Pr(\hat{y}_{i}\neq y_{i}\mid\bm{\beta}^{*})=\sum\limits_{i=1}^{n}\Pr(\hat{y}_{i}\neq y_{i}\mid\bm{\beta}^{*}_{\mathcal{A}})=\text{err}(\hat{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})$, where $\hat{y}_{i}\sim\text{Binomial}(1,\frac{1}{1+\exp(-\hat{\bm{\beta}}_{\mathcal{A}}^{\top}{\bm{x}_{i}}_{\mathcal{A}})})$. Similarly, the following conclusion also holds true, $\text{err}(\tilde{\bm{\beta}};\bm{\beta}^{*})=\text{err}(\tilde{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})$. Before proving the theorem, a lemma about the asymptotic excess errors is needed.
@@ A-B Proof of Throrem 2
$\lim_{n\to\infty}n\left[\mathbb{E}\left(\text{err}(\hat{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right)-\text{err}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right]$
$=$ $\operatorname{trace}\{J(\bm{\beta}^{*}_{\mathcal{A}})I^{-1}_{\mathcal{A}}\}$ (49)
@@
$\lim_{n\to\infty}n\left[\mathbb{E}\left(\text{err}(\tilde{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right)-\text{err}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right]$
$=$ $\frac{m+\alpha_{0}}{m(1+\alpha_{0})}\operatorname{trace}\{J(\bm{\beta}^{*}_{\mathcal{A}})I^{-1}_{\mathcal{A}}\},$ (50)
@@
Proof:
First, consider the Taylor expansion of the error function $\text{err}(\hat{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})$ around the true parameter $\bm{\beta}^{*}_{\mathcal{A}}$:
$\text{err}(\hat{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})$ $=\text{err}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})+(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}})^{\top}\text{err}^{{}^{\prime}}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})$
$+\frac{(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}})^{\top}\text{err}^{{}^{\prime\prime}}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}})}{2}$
$+o(\|\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}}\|_{2}^{2})$
@@
$\lim_{n\to\infty}n\left[\mathbb{E}\left(\text{err}(\hat{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right)-\text{err}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right]$
$=$ $\lim_{n\to\infty}n\mathbb{E}\left[\frac{(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}})^{\top}\text{err}^{{}^{\prime\prime}}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}})}{2}\right].$
@@
$=$ $\operatorname{trace}\left(\text{err}^{{}^{\prime\prime}}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\cdot\text{Cov}(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}})\right)$
$+$ $\mathbb{E}(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}})^{\top}\text{err}^{{}^{\prime\prime}}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\mathbb{E}(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}}).$
@@
$\lim_{n\to\infty}n\mathbb{E}\left[\frac{(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}})^{\top}\text{err}^{{}^{\prime\prime}}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}})}{2}\right]$
$=$ $\frac{1}{2}\operatorname{trace}\left(\text{err}^{{}^{\prime\prime}}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\cdot I^{-1}_{\mathcal{A}}\right).$
@@
By implementing the same steps on
$\lim_{n\to\infty}n\left[\mathbb{E}\left(\text{err}(\tilde{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right)-\text{err}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right]$, it can be concluded that
$\lim_{n\to\infty}n\left[\mathbb{E}\left(\text{err}(\tilde{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right)-\text{err}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right]$
$=$ $\frac{m+\alpha_{0}}{m(1+\alpha_{0})}\operatorname{trace}\{J(\bm{\beta}^{*}_{\mathcal{A}})I^{-1}_{\mathcal{A}}\}.$
This completes the proof of the lemma. ∎
@@
A-C Proof of Throrem 3
A-C1 Basic Lemma and Four Matrices

### v2

@@ 5.2 Real Data Experiments
In this section, we validate that manual labeling can enhance the variable selection ability of PLR on datasets in Table 3. These datasets is available at https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/. Due to space limitations, we only present the numerical results of the w8a dataset with the largest data size in the table. The numerical results of other datasets can be found in supplementary material 1.5.
Table 3: Basic information of datasets.
ijcmn1 phishing a9a w8a
Features (d) 22 68 123 300
Samples (n) 49990 11055 32561 49749
@@
We randomly partitioned the original dataset comprising 49,749 observations into a training set of 40,000 observations and a test set of 9,749 observations. Since our algorithm exhibits partition insensitivity, to save computation time and account for the device’s memory, we set $G=20$ in the parallel computing environment. Table 4 presents the simulated relative efficiency of PLR. The true coefficient $\bm{\beta}^{*}$, which serves to generate a posterior probability for manual labeling, is unknown. Here, it can be estimated using the entire sample data. The simulated ARE values tend to increase as $\alpha_{0}$ and $m$ increase, which is consistent with the conclusion of Theorem 2. It is observed that the simulated relative efficiencies are much lower than their corresponding theoretical values. This discrepancy might stem from two factors. Firstly, there could be an underestimation of the conditional error rate $\text{err}(\hat{\bm{\beta}};\bm{\beta}^{*})$ in (3.4). Secondly, the true conditional error rate $\text{err}({\bm{\beta}}^{*};\bm{\beta}^{*})$ may not be precisely estimated using the limited test set. In the simulations presented in Section 5.1, we are able to compute the exact $\text{err}(\hat{\bm{\beta}};\bm{\beta}^{*})$ and $\text{err}({\bm{\beta}}^{*};\bm{\beta}^{*})$. This explains why the estimated ARE is closer to the theoretical ARE in Section 5.1.
@@
Table 4: The asymptotic relative efficiency (ARE) on the w8a dataset. The theoretical ARE of PLR, as stated in Theorem 2, is provided in parentheses.
@@
6 Conclusion and Future Work
This paper theoretically proves that when label noise is only related to classification difficulty, it is useful for variable selection in binary classification PLR. This benefit will be reflected in more accurate estimates of important coefficients compared to using truth labels. In addition, we also propose a new parallel algorithm for solving the large-scale PLR with manually generated labels. This parallel algorithm is insensitive to data partitioning, meaning that the solution of the algorithm remains unchanged regardless of different data partitions.
However, the theoretical results in this paper are only applicable when $p<n$, that is, when the feature dimension is smaller than the number of samples. Extending the beneficial noise concept to high-dimensional cases ($p>n$) or even ultra-high-dimensional cases ($p\gg n$) is a challenging task. Future research can be extended in several directions: expanding to multi-class classifier [37] and multi-view classifier [38] to see if the beneficial effect of label noise persists; optimizing the parallel algorithm for better computational efficiency and scalability using techniques like GPU acceleration; incorporating temporal and spatial information into the PLR model and algorithm to handle data with such characteristics.
Appendix A Proof of Throrem 1
@@
$\Gamma_{n}(\mathbf{u})=\sum_{i=1}^{n}\left(-S_{i}\left(\bm{x}_{i}^{\top}\left(\bm{\beta}^{*}+\frac{\bm{u}}{\sqrt{n}}\right)\right)+\phi\left(\bm{x}_{i}^{\top}\left(\bm{\beta}^{*}+\frac{\bm{u}}{\sqrt{n}}\right)\right)\right)+\lambda\sum_{j=1}^{d}w_{j}\left|\beta_{j}^{*}+\frac{u_{j}}{\sqrt{n}}\right|.$
Let $\tilde{\bm{u}}^{(n)}=\arg\min\limits_{\bm{u}}\Gamma_{n}(\bm{u})$; then $\tilde{\bm{u}}^{(n)}=\sqrt{n}(\tilde{\bm{\beta}}-\bm{\beta}^{*})$, where $\tilde{\bm{\beta}}$ is the PLR estimator with manual labels in (3.2).
@@
$\mathbb{E}_{S_{i},\bm{x}_{i}}\left(\left[S_{i}-\phi^{\prime}\left(\bm{x}_{i}^{\top}\bm{\beta}^{*}\right)\right]\bm{x}_{i}^{\top}\right)=\bm{0}$ (S1.1)
@@
$\text{Var}_{S_{i},\bm{x}_{i}}\left(\left[S_{i}-\phi^{\prime}\left(\bm{x}_{i}^{\top}\bm{\beta}^{*}\right)\right]\bm{x}_{i}^{\top}\right)=\frac{m+\alpha_{0}}{1+\alpha_{0}}mI(\bm{\beta}^{*}).$ (S1.2)
@@
$\text{Var}_{S_{i},\bm{x}_{i}}\left(\left[S_{i}-\phi^{\prime}\left(\bm{x}_{i}^{\top}\bm{\beta}^{*}\right)\right]\left(\bm{x}_{i}^{\top}\bm{u}\right)\right)=\bm{u}^{\top}\frac{m+\alpha_{0}}{1+\alpha_{0}}mI(\bm{\beta}^{*})\bm{u}.$ (S1.3)
@@
$A_{1}^{(n)}\xrightarrow{d}\bm{u}^{\top}N(0,\frac{m+\alpha_{0}}{1+\alpha_{0}}mI(\bm{\beta}^{*})).$ (S1.4)
For the second term $A_{2}^{(n)}$, from (2.2), we observe that
$\sum_{i=1}^{n}\phi^{\prime\prime}\left(\bm{x}_{i}^{\top}\bm{\beta}^{*}\right)\frac{\left(\bm{x}_{i}\bm{x}_{i}^{\top}\right)}{n}\xrightarrow{p}mI(\bm{\beta}^{*}),$ (S1.5)
where $\xrightarrow{p}$ represents convergence in probability. Thus,
$A_{2}^{(n)}\xrightarrow{p}\frac{m}{2}\bm{u}^{\top}I(\bm{\beta}^{*})\bm{u}.$ (S1.6)
@@
$\frac{\lambda}{\sqrt{n}}{w}_{j}\sqrt{n}\left(\left|\beta_{j}^{*}+\frac{u_{j}}{\sqrt{n}}\right|-\left|\beta_{j}^{*}\right|\right)\xrightarrow{p}\begin{cases}0&\text{if }\beta_{j}^{*}\neq 0\\ 0&\text{if }\beta_{j}^{*}=0\text{ and }u_{j}=0\\ \infty&\text{if }\beta_{j}^{*}=0\text{ and }u_{j}\neq 0.\end{cases}$ (S1.7)
@@
$6\sqrt{n}A_{4}^{(n)}\leq$ $\sum_{i=1}^{n}\frac{1}{n}M(\bm{x}_{i}^{\top})|\bm{x}_{i}^{\top}\bm{u}|^{3}\xrightarrow{p}E[M(\bm{x})|\bm{x}^{\top}\bm{u}|^{3}]<\infty.$ (S1.8)
Thus, based on the results from (S1.4), (S1.6), (S1.7) and (S1.8), by invoking Slutsky’s theorem, we can infer that $H^{(n)}(\bm{u})\xrightarrow{d}H(\bm{u})$ for every $\bm{u}$, where
@@
$\tilde{\bm{u}}_{\mathcal{A}}^{(n)}\xrightarrow{d}N\{0,\frac{m+\alpha_{0}}{m(1+\alpha_{0})}I_{\mathcal{A}}^{-1}\}\quad\text{and}\quad\tilde{\bm{u}}_{\mathcal{A}^{c}}^{(n)}\xrightarrow{d}\bm{0}.$ (S1.9)
@@
$\sum_{i=1}^{n}x_{ij^{\prime}}\left(S_{i}-\phi^{\prime}\left(\bm{x}_{i}^{\top}\tilde{\bm{\beta}}\right)\right)=\lambda w_{j};$ (S1.10)
@@
where $\tilde{\bm{\beta}}^{*}$ is between $\tilde{\bm{\beta}}$ and $\bm{\beta}$$^{*}$. From (S1.1) and (S1.2), we know that
$B_{1}^{(n)}\xrightarrow{d}N(0,\frac{m+\alpha_{0}}{1+\alpha_{0}}mI(\bm{\beta}^{*})).$
It follows from (S1.5) that
$\frac{1}{n}\sum_{i=1}^{n}x_{ij^{\prime}}\phi^{\prime\prime}\left(\bm{x}_{i}^{\top}{\bm{\beta}}^{*}\right)\times\bm{x}_{i}^{\top}\xrightarrow{p}mI_{j^{{}^{\prime}}}(\bm{\beta}^{*}),$
where $I_{j^{{}^{\prime}}}(\bm{\beta}^{*})$ is the $j^{{}^{\prime}}$-th row of $I(\bm{\beta}^{*})$. Combining (S1.9), it can be concluded that $B_{2}^{(n)}$ converges to some normal random variable. It follows from the regularity condition (A2) and (S1.9) that $B_{3}^{(n)}=O_{p}(1/\sqrt{n})$. Meanwhile, since $\bm{\beta}^{{}^{\prime}}$ is a $\sqrt{n}$-consistent estimator of $\bm{\beta}^{*}$, we have $\sqrt{n}{\beta}_{j}^{\prime}\xrightarrow{p}O(1)$. Then,
$\frac{\lambda w_{j}}{\sqrt{n}}=\frac{\lambda}{\sqrt{n}}n^{\gamma/2}\frac{1}{\left|\sqrt{n}{\beta}_{j}^{\prime}\right|^{\gamma}}\xrightarrow{p}\infty.$
By considering the left and right sides of the (S1.10), it can be concluded that $\Pr(j^{\prime}\in\tilde{\mathcal{A}})\to 0$. This completes the proof of Theorem 1.
Appendix B Proof of Throrem 2
From the conclusions of Section 2.1.2 and Theorem 1, it can be known that when $n\to\infty$, $\bm{\beta}^{*}=({\bm{\beta}^{*}_{\mathcal{A}}}^{\top},\bm{0}^{\top})^{\top}$, $\hat{\bm{\beta}}=(\hat{\bm{\beta}}_{\mathcal{A}}^{\top},\bm{0}^{\top})^{\top}$ and $\tilde{\bm{\beta}}=(\tilde{\bm{\beta}}_{\mathcal{A}}^{\top},\bm{0}^{\top})^{\top}$. This also means that $\text{err}(\hat{\bm{\beta}};\bm{\beta}^{*})=\sum\limits_{i=1}^{n}\Pr(\hat{y}_{i}\neq y_{i}\mid\bm{\beta}^{*})=\sum\limits_{i=1}^{n}\Pr(\hat{y}_{i}\neq y_{i}\mid\bm{\beta}^{*}_{\mathcal{A}})=\text{err}(\hat{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})$, where $\hat{y}_{i}\sim\text{Binomial}(1,\frac{1}{1+\exp(-\hat{\bm{\beta}}_{\mathcal{A}}^{\top}{\bm{x}_{i}}_{\mathcal{A}})})$. Similarly, the following conclusion also holds true, $\text{err}(\tilde{\bm{\beta}};\bm{\beta}^{*})=\text{err}(\tilde{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})$. Before proving the theorem, a lemma about the asymptotic excess errors is needed.
@@ Appendix B Proof of Throrem 2
$\lim_{n\to\infty}n\left[\mathbb{E}\left(\text{err}(\hat{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right)-\text{err}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right]=\operatorname{trace}\{J(\bm{\beta}^{*}_{\mathcal{A}})I^{-1}_{\mathcal{A}}\}$
@@
$\lim_{n\to\infty}n\left[\mathbb{E}\left(\text{err}(\tilde{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right)-\text{err}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right]=\frac{m+\alpha_{0}}{m(1+\alpha_{0})}\operatorname{trace}\{J(\bm{\beta}^{*}_{\mathcal{A}})I^{-1}_{\mathcal{A}}\},$
@@
Proof. First, consider the Taylor expansion of the error function $\text{err}(\hat{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})$ around the true parameter $\bm{\beta}^{*}_{\mathcal{A}}$:
$\text{err}(\hat{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})$
[…]
@@
$\lim_{n\to\infty}n\left[\mathbb{E}\left(\text{err}(\hat{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right)-\text{err}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right]=\lim_{n\to\infty}n\mathbb{E}\left[\frac{(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}})^{\top}\text{err}^{{}^{\prime\prime}}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}})}{2}\right].$
@@
[…]
@@
$\lim_{n\to\infty}n\mathbb{E}\left[\frac{(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}})^{\top}\text{err}^{{}^{\prime\prime}}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})(\hat{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}})}{2}\right]=\frac{1}{2}\operatorname{trace}\left(\text{err}^{{}^{\prime\prime}}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\cdot I^{-1}_{\mathcal{A}}\right).$
@@
By implementing the same steps on $\lim_{n\to\infty}n\left[\mathbb{E}\left(\text{err}(\tilde{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right)-\text{err}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right]$, it can be concluded that
$\lim_{n\to\infty}n\left[\mathbb{E}\left(\text{err}(\tilde{\bm{\beta}}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right)-\text{err}(\bm{\beta}^{*}_{\mathcal{A}};\bm{\beta}^{*}_{\mathcal{A}})\right]=\frac{m+\alpha_{0}}{m(1+\alpha_{0})}\operatorname{trace}\{J(\bm{\beta}^{*}_{\mathcal{A}})I^{-1}_{\mathcal{A}}\}.$
This completes the proof of the lemma.
@@
Appendix C Proof of Throrem 3
C.1 Basic Lemma and Four Matrices